Grade 11Mechanics

n identical cubes each of mass m and edge L are on a floor .if the cubes are to be arranged one over the other in a vertical stack ,the work to be done is

sadvik reddy
8 Years agoGrade 11

2 Answers

Arun
8 Years ago
Work done in placing a 2nd cube  on the first one, W2​ mgl 
Similarly  W3 = 2 mgl                                                                 
                W4 = 3 mgl 
               Wn = (n -1) mgl 
Total work done  WN = W1+ W2+ W3...+Wn
WN​ = [ mgl+2mgl+3mgl....(n - 1) mgl ]
WN​ =mgl [ 1+2+3+......(n - 1) ]
Chintan Shah
7 Years ago
Work done in placing first cube is 0.
Work done in placing 2nd cube is mgl.
Similarly, W3=2mgl,.....,Wn=(n-1)mgl.
 Wnet= W1+W2+W3+......+W(n-1)
          = mgl{1+2+3+....+(n-1)}
          = mgl{Sn}
          = mgl{(n-1)/2 [2×1+(n-1-1)1]}
          = mgl{(n-1)/2×n}
Wnet = mgln(n-1)/2.